OpenStudy (anonymous):

Find the sum of the coordinates of the solution of: 17x + 31y = 274 7x − 5y = 24 (A) 6 (B) 8 (C) 10 (D) 12

OpenStudy (anonymous):

solve for x or y in one of the equations first

OpenStudy (anonymous):

7x-5y=24 5y=7x-24 y=7/5x - 24/5 = 1.4x - 4.8 now substitute the y value into the second equation and solve for x.

OpenStudy (anonymous):

17x + 31(1.4x - 4.8)=274 17x + 43.4x - 148.8=274 60.4x=274+148.8 60.4x=422.8 x=7 now solve for y using y=1.4x - 4.8

OpenStudy (anonymous):

y=1.4*7-4.8=5 the sum of coordinates will be equal to x+y=7+5=12

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

np :)